In an unhealthy, dusty cement mill, there was 2.6x10




9 dust particles (sp g =3.0) per




cubic meter of air. Assuming the particles to be spheres of 2.0 µm diameter, calculate




the mass of dust (a) in a 20mx15mx80m room and (b) inhaled in each average




breath of 400-cm3 volume

1
Expert's answer
2021-12-14T13:33:31-0500

"m=N\rho V_0,"

"N=nV,"

"V_0=\frac{\pi d^3}6,"

"m=\frac{n\rho\pi d^3}6 V,"

a)

"m=784~g,"

b)

"m=13~ng."